# Tracking Transforming Objects: A Benchmark

### v1

@@ Abstract
Tracking transforming objects holds significant importance in various fields due to the dynamic nature of many real-world scenarios. By enabling systems accurately represent transforming objects over time, tracking transforming objects facilitates advancements in areas such as autonomous systems, human-computer interaction, and security applications. Moreover, understanding the behavior of transforming objects provides valuable insights into complex interactions or processes, contributing to the development of intelligent systems capable of robust and adaptive perception in dynamic environments. However, current research in the field mainly focuses on tracking generic objects. In this study, we bridge this gap by collecting a novel dedicated Dataset for Tracking Transforming Objects, called DTTO, which contains 100 sequences, amounting to approximately 9.3K frames. We provide carefully hand-annotated bounding boxes for each frame within these sequences, making DTTO the pioneering benchmark dedicated to tracking transforming objects. We thoroughly evaluate 20 state-of-the-art trackers on the benchmark, aiming to comprehend the performance of existing methods and provide a comparison for future research on DTTO. With the release of DTTO, our goal is to facilitate further research and applications related to tracking transforming objects.
@@
00footnotetext: These authors contributed equally to this work.Object tracking stands as one of the cornerstone challenges in computer vision, finding application across diverse fields such as video surveillance, robotics, human-machine interaction, and motion analysis [1, 2, 3, 4, 5, 6]. The advent of deep learning has notably propelled significant advancements in object tracking in recent years [7, 8, 9, 10, 11, 12, 13, 14]. However, despite substantial progress in the field, current research predominantly concentrates on tracking generic objects, with limited attention directed towards tracking transforming objects. Although exploratory studies have been conducted on segmenting transforming objects [15, 16, 17] in video object segmentation tasks very recently, their objective is primarily to predict pixel-wise masks of objects undergoing transformations in each frame of the video, utilizing the initial frame mask as a reference. Moreover, these studies maintain the object category consistent throughout each sequence. In this work, we investigate the tracking of transforming objects, which not only undergo apparent changes in appearance, context, or viewpoint but also may change categories during the transformation process. This presents a more conceptually challenging task due to the involvement of more complex cognitive processes [18, 19, 20, 21]. For instance, humans possess the ability to maintain the consistent identity of objects despite alterations in appearance, context, or viewpoint. This cognitive ability, known as object constancy, enables individuals to perceive an object’s underlying identity regardless of variations in its form or visual representation [22, 19]. Moreover, humans demonstrate cognitive flexibility, allowing them to adjust their perceptions and comprehension based on new information or experiences [18, 23, 21]. This adaptability enables individuals to recognize an object’s identity even as it undergoes changes in form, such as witnessing a robot transform into a car within the context of a movie. These cognitive processes are essential for humans to successfully track transforming objects, yet none of the mainstream discriminative and generative methods take these into account. This study introduces the task of tracking transforming objects by predicting the bounding box for the transforming object in each frame of the video, which explicitly involves these cognitive processes and present more conceptual challenges to tracking algorithms.
@@ 3.1 Data Collection
Our benchmark aims to cover common object classes that undergo transformations during tracking. To achieve this goal, we select 11 categories, including ‘building’, ‘car’, ‘cell’, ‘cactus’, ‘container’, ‘flower’, ‘frog’, ‘mushroom’, ‘seed’, ‘toy’, and ‘transformer’, to develop the DTTO. Among them, we present six types of transformations: ‘bloom’, ‘build’, ‘collide’, ‘crush’, ‘grow’, and ‘transform’. After identifying the object categories and transformation types, we collected a variety of video sequences from the Internet, as it serves as a primary source for many tracking benchmarks (such as LaSOT [25], GOT-10k [24], TrackingNet [26], etc.). At first, we collected more than 150 sequences, totaling over 30K frames. While all the previously selected sequences exhibit transforming objects, not all lead to a significant transformation. Therefore, we eliminate irrelevant sequences and redundant frames to ensure that the objects presented in the remaining sequences undergo a substantial transformation process. Ultimately, the DOTT dataset comprises 100 video sequences, totaling 9,297 frames, with an average video length of approximately 93 frames. Fig. 2 presents the histogram depicting the distribution of transformation types across video sequences (left) and co-occurrence statistics between transformation types and object categories (right), respectively. As evident, the ’transform’, ’grow’, and ’bloom’ transformations are the most common, collectively constituting over 75% of all sequences. It’s noteworthy that these transformations involve changes in object categories, underscoring the conceptual complexities inherent in our dataset. In addition, there is significant entropy in the correlation statistics, highlighting the diversity of our dataset.
@@ 4.3 Evaluation Results
(a) Attribute-based comparison of SV, ARC, and ROT in terms of precision.
(b) Attribute-based comparison of SV, ARC, and ROT in terms of success rate.
Attribute-based performance. To gain a better analysis of the differences in performance across various trackers, we evaluated their performance across eleven attributes. Due to page limitations, Fig. 6 displays the attribute-based evaluation results for the three most prevalent attributes: SV, ARC, and ROT. As illustrated in Fig. 6 (a), GRM demonstrates the highest precision, i.e., 73.0%, 75.2%, and 74.5%, across all three attribute subsets, with DropTrack and OSTrack securing the second position. Specifically, DropTrack ranks second in ARC and SV with precisions of 72.2% and 72.5%, while OSTrack ranks second in ROT with 72.7% precision. In terms of success rates, as shown in Fig. 6 (b), GRM and DropTrack occupy the top two positions in ASC and SV. Specifically, their SUCs are 74.3% and 74.2% for ASC, and 73.3% and 73.1% for SV, respectively. Additionally, GRM and ROMTrack achieve the best two results in ROT with 77.7% and 76.5%, respectively. These results suggest that GRM and DropTrack are basically the top-performing trackers on these attribute subsets.

### v2

@@ Abstract
Tracking transforming objects holds significant importance in various fields due to the dynamic nature of many real-world scenarios. By enabling systems accurately represent transforming objects over time, tracking transforming objects facilitates advancements in areas such as autonomous systems, human-computer interaction, and security applications. Moreover, understanding the behavior of transforming objects provides valuable insights into complex interactions or processes, contributing to the development of intelligent systems capable of robust and adaptive perception in dynamic environments. However, current research in the field mainly focuses on tracking generic objects. In this study, we bridge this gap by collecting a novel dedicated Dataset for Tracking Transforming Objects, called DTTO, which contains 100 sequences, amounting to approximately 9.3K frames. We provide carefully hand-annotated bounding boxes for each frame within these sequences, making DTTO the pioneering benchmark dedicated to tracking transforming objects. We thoroughly evaluate 20 state-of-the-art trackers on the benchmark, aiming to comprehend the performance of existing methods and provide a comparison for future research on DTTO. With the release of DTTO, our goal is to facilitate further research and applications related to tracking transforming objects. The dataset and evaluation toolkit are available at https://github.com/wuyou3474/DTTO.
@@
00footnotetext: These authors contributed equally to this work.Thanks to the support by Guangxi Science and Technology Base and Talent Special Project (No. Guike AD22035127).
Object tracking stands as one of the cornerstone challenges in computer vision, finding application across diverse fields such as video surveillance, robotics, human-machine interaction, and motion analysis [1, 2, 3, 4, 5, 6, 7, 8, 9]. The advent of deep learning has notably propelled significant advancements in object tracking in recent years [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. However, despite substantial progress in the field, current research predominantly concentrates on tracking generic objects, with limited attention directed towards tracking transforming objects. Although exploratory studies have been conducted on segmenting transforming objects [20, 21, 22] in video object segmentation tasks very recently, their objective is primarily to predict pixel-wise masks of objects undergoing transformations in each frame of the video, utilizing the initial frame mask as a reference. Moreover, these studies maintain the object category consistent throughout each sequence. In this work, we investigate the tracking of transforming objects, which not only undergo apparent changes in appearance, context, or viewpoint but also may change categories during the transformation process. This presents a more conceptually challenging task due to the involvement of more complex cognitive processes [23, 24, 25]. For instance, humans possess the ability to maintain the consistent identity of objects despite alterations in appearance, context, or viewpoint. This cognitive ability, known as object constancy, enables individuals to perceive an object’s underlying identity regardless of variations in its form or visual representation [26, 24]. Moreover, humans demonstrate cognitive flexibility, allowing them to adjust their perceptions and comprehension based on new information or experiences [23, 27, 25]. This adaptability enables individuals to recognize an object’s identity even as it undergoes changes in form, such as witnessing a robot transform into a car within the context of a movie. These cognitive processes are essential for humans to successfully track transforming objects, yet none of the mainstream discriminative and generative methods take these into account. This study introduces the task of tracking transforming objects by predicting the bounding box for the transforming object in each frame of the video, which explicitly involves these cognitive processes and presents more conceptual challenges to tracking algorithms.
@@ 3.1 Data Collection
Our benchmark aims to cover common object classes that undergo transformations during tracking. To achieve this goal, we select 11 categories, including ‘building’, ‘car’, ‘cell’, ‘cactus’, ‘container’, ‘flower’, ‘frog’, ‘mushroom’, ‘seed’, ‘toy’, and ‘transformer’, to develop the DTTO. Among them, we present six types of transformations: ‘bloom’, ‘build’, ‘collide’, ‘crush’, ‘grow’, and ‘transform’. After identifying the object categories and transformation types, we collected a variety of video sequences from the Internet, as it serves as a primary source for many tracking benchmarks (such as LaSOT [29], GOT-10k [28], TrackingNet [30], etc.). At first, we collected more than 150 sequences, totaling over 30K frames. While all the previously selected sequences exhibit transforming objects, not all lead to a significant transformation. Therefore, we eliminate irrelevant sequences and redundant frames to ensure that the objects presented in the remaining sequences undergo a substantial transformation process. Ultimately, the DTTO dataset comprises 100 video sequences, totaling 9,297 frames, with an average video length of approximately 93 frames. Fig. 2 presents the histogram depicting the distribution of transformation types across video sequences (left) and co-occurrence statistics between transformation types and object categories (right), respectively. As evident, the ’transform’, ’grow’, and ’bloom’ transformations are the most common, collectively constituting over 75% of all sequences. It’s noteworthy that these transformations involve changes in object categories, underscoring the conceptual complexities inherent in our dataset. In addition, there is significant entropy in the correlation statistics, highlighting the diversity of our dataset.
@@ 4.3 Evaluation Results
Attribute-based performance. To gain a better analysis of the differences in performance across various trackers, we evaluated their performance across eleven attributes. Due to page limitations, Fig. 6 displays the attribute-based evaluation results for the three most prevalent attributes: SV, ARC, and ROT. As illustrated in Fig. 6 (a), GRM demonstrates the highest precision, i.e., 74.5%, 73.0%, and 75.2%, across all three attribute subsets, with DropTrack and OSTrack securing the second position. Specifically, DropTrack ranks second in ARC and SV with precisions of 72.2% and 72.5%, while OSTrack ranks second in ROT with 72.7% precision. In terms of success rate, as shown in Fig. 6 (b), GRM and DropTrack occupy the top two positions in ASC and SV. Specifically, their SUCs are 74.3% and 74.2% for ASC, and 73.3% and 73.1% for SV, respectively. Additionally, GRM and ROMTrack achieve the best two results in ROT with 77.7% and 76.5%, respectively. These results suggest that GRM and DropTrack are basically the top-performing trackers on these attribute subsets.